OpenStudy (anonymous):

Rita traveled 1,250 miles in the first 3 days of her trip. At this rate, how long will it take her to travel 1,875 miles?

OpenStudy (anonymous):

medal will go to...

OpenStudy (esshotwired):

Use this equation. 1875(3) = 1250x It is pretty much like a ratio.

OpenStudy (anonymous):

esshotwired of course :)

OpenStudy (esshotwired):

1875(3)=1250x 5625=1250x 4.5=x So it would take her 4.5 days to travel 1875 miles.

OpenStudy (anonymous):

X = no. of days 1875/1250 = X/3 1.5 * 3 = X X = 4.5 days = 5 days since 4.5 can't be a day unless it's allowed if the question is asking how many days left then it would be 5 - 3 days = 2 days

OpenStudy (anonymous):

@esshotwired im confused on the second step how did u get 4.5?

OpenStudy (esshotwired):

you have to cancel out 1250 so you have just x so you divide 1250 from both sides

OpenStudy (anonymous):

OHH I GET IT THANKS

OpenStudy (esshotwired):

Your welcome :)